OpenStudy (anonymous):

Use logarithms to solve the equation 8e^(3x) + 4 = 15 Round to the nearest thousandth. Explain you answer step by step please

OpenStudy (anonymous):

1) subtract 4 2) divide by 8 3) write in equivalent logarithmic form 4) divide by 3

OpenStudy (anonymous):

it would be 0.106 ?

OpenStudy (anonymous):

i don't know did you compute \[\frac{\ln(\frac{11}{8})}{3}\]?

OpenStudy (anonymous):

yeah that looks good to me

OpenStudy (queelius):

Further "simplifications" that can be done: \[\frac { \ln { \frac{ 11 }{ 8 } } } { 3 } = \ln { (\frac{ 11 }{ 8 })^{1/3} } = \ln { (\frac{ 11^{1/3} }{ 2 }) } = \frac { 1 }{ 3} \ln(11) - \ln(2)\]